A ball is shot from the ground into the air.At a height of 9.1m ,its velocity is found to be 7.6i+6.1j m/s. To what maximum height the ball will raise. 

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear sindhuja

velocity at hight 9.1m is 7.6i + 6.1j m/s

it it wil raise hight h

u=6.1 m/s (verticle component)

v=0

v2=u2-2gh

0=6.12-2g h

h=6.12/2g

h=1.9 m

so total hight =9.1+1.9 =11m


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin